Four of the angles of a pentagon have measures 30, 70, 120, and 165. Find the measure of the fifth angle. Please show how it is done.
Answer by checkley75(3666)

THE SUM OF THE ANGLS OF A PENTAGON ARE 90(5+1)=540
30+70+120+165=385
540-385=155 DEGREES FOR THE LAST ANGLE.